OpenStudy (anonymous):

x= 1+ √ ̅x+5 x+5 is under one radical the two answers are 4 and -1, but the teacher said -1 rejects; however, when i plug in the numbers, it works out. can someone explain why -1 rejects please?

OpenStudy (mertsj):

-1 IS NOT equal to 1 + sqrt4

OpenStudy (mertsj):

So -1 does not make the sentence true and must be rejected.

OpenStudy (ranga):

x = 1 + sqrt(x+5) (x-1) = sqrt(x+5) square both sides: (x-1)^2 = x+5 x^2 - 2x + 1 - x - 5 = 0 x^2 - 3x - 4 = 0 (x - 4)(x + 1) = 0 x = 4 or x = -1 Try 4 in the original equation: x = 1 + sqrt(x+5) LHS = 4 RHS = 1 + sqrt(4 + 5) = 1 + sqrt(9) = 1 + 3 = 4 LHS = RHS. x = 4 is a valid solution. Try -1 LHS = -1 RHS = 1 + sqrt(-1+5) = 1 + sqrt(4) = 1 + 2 = 3 LHS is not equal to RHS. x = -1 is NOT a solution.

OpenStudy (anonymous):

oh i kinda see. what does lhs and rhs stand for?

OpenStudy (mertsj):

left hand side and right hand side

OpenStudy (anonymous):

ohh so u have to see if x-1 = √x+5 ?

OpenStudy (anonymous):

ok i understand now. thanks so much!!!!